Blood flow is partially controlled by the cross-sectional area of the blood vessel through which the blood is traveling. Blood vessel B has a radius of 1 times 10−4mm and blood vessel D has a radius of 5 times 10−5mm. Let r1 be the radius of B and r2 be the radius of D. Find the ratio of r1 to r2. What does the ratio tell you?
Answer by stanbon(75887)

Blood flow is partially controlled by the cross-sectional area of the blood vessel through which the blood is traveling. Blood vessel B has a radius of 1 times 10−4mm and blood vessel D has a radius of 5 times 10−5mm. Let r1 be the radius of B and r2 be the radius of D. Find the ratio of r1 to r2. What does the ratio tell you?
------
r1:r2 = r1/r2 = (1x10^(-4))/(5x10^-5)
= (1/5)x10
= 0.2x10
= 2
---------------
r1:r2 = 2:1
----------------
r1 is twice the radius of r2
===============================
Cheers,
Stan H.